OpenStudy (anonymous):

5. What is the solution to the system of equations shown below? -2x + y = 10 2x + 4y = -30 (Points : 2) (3, 16) (2, -7) (-7, -4) (-2, 7)

OpenStudy (anonymous):

-2x + y = 10 2x + 4y = -30 add them 0 + 5y = -20 divide 5 to both sides y = -4 to find the x, plug the y value in one of the equation and solve for x